OpenStudy (anonymous):

Solve the given polynomial equation. Use the ratinal zero theorm, descrate's rule of signs, and possibly the graph of the polynomial function shown by graphing utility as an aid in obtaiing the first root. 7x^5+41x^4+43x^3-91x^2-114x+18=0 find the soloution set?

OpenStudy (anonymous):

\[(x+3)^2 (7 x-1) \left(x^2-2\right)=0 \]

OpenStudy (anonymous):

so thats the solution?

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

\[x=\left\{-3,-3,\frac{1}{7},-\sqrt{2},\sqrt{2}\right\} \]

OpenStudy (anonymous):

x^4+3x^3+3x^2+3x+2=0 so this one whoudl be the same?